# A new class of positive linear operators preserving logarithmic functions

### v1

@@ Abstract
In this paper, we introduce a new class of positive linear operators that generalize the classical Bernstein operators. Specifically, we construct a sequence of operators that reproduce the logarithmic function $\ln(1+\mu+x)$, with $\mu>0$ and $x\in[0,1]$. We prove pointwise and uniform convergence and we derive a quantitative estimate of the approximation error in terms of the modulus of continuity. We also obtain a Voronovskaja-type asymptotic formula, that is used to establish saturation results and inverse theorems. In particular, the saturation class of the considered approximation process is characterized by solving a second order differential equation. Shape-preserving properties, such as monotonicity, concavity and variation diminishing, are also investigated. Finally, a simple application to signal denoising is addressed.
@@ 1 Introduction
Over the years, many extensions of the Bernstein polynomials have been proposed in the literature, including, for example, the sampling-type operators ([12, 13]), the max-product operators ([14]), the neural network-type operators ([25]), the Szasz-Mirakjan operators ([1, 2, 23]), and many others ([22, 3, 30, 5, 21, 20, 24]).
@@
with $x\in[0,1]$. Because of their connection with the classical operators $B_{n}$, it can be shown that $\mathscr{G}_{n}$ exactly reproduces the exponential functions $\exp_{\mu}(x):=e^{\mu x}$ and $\exp_{\mu}^{2}(x):=e^{2\mu x}$, i.e., for $x\in[0,1]$,
@@
Within this line of research, in the present paper we introduce a new class of positive linear operators, below denoted by $\mathscr{L}_{n}$, that preserve the logarithmic function $\ln_{\mu}(x)=\ln(1+\mu+x)$, with $\mu>0$ and $x\in[0,1]$. The idea is to replace the exponential weights in $\mathscr{G}_{n}$ with suitable logarithmic ones. Moreover, the functions $\tilde{a}_{n}$ are replaced by suitable concave approximants of the identity function involving $\ln_{\mu}(x)$. Differently from the exponential type operators, in the literature there is a lack of results concerning operators that preserve logarithmic functions. Besides this, a further motivation for the introduction of the operators $\mathscr{L}_{n}$ is the fact that the properties of the logarithmic function could be useful to linearize nonlinear transformations of multiplicative type.
@@
2 Preliminaries
Let us provide some notations, definitions and classical results that will be used throughout the paper.
We recall that the binomial weights
$p_{n,k}(x):=\binom{n}{k}x^{k}(1-x)^{n-k},\quad x\in[0,1],\quad k=0,\dots,n,\quad n\in\mathbb{N},$
play a fundamental role in the study of approximation processes. The integrals of $p_{n,k}(x)$ can be computed explicitly using the Euler $\Gamma$ function, as follows:
$\displaystyle\int_{0}^{1}p_{n,k}(x)dx$ $=\binom{n}{k}\int_{0}^{1}x^{k}(1-x)^{n-k}dx$
$=\binom{n}{k}\frac{\Gamma(k+1)\Gamma(n-k+1)}{\Gamma(n+2)}$
$=\binom{n}{k}\frac{k!(n-k)!}{(n+1)!}=\frac{1}{n+1},$ (1)
where $\Gamma(k)=(k-1)!$ for every positive integer $k$. Moreover, for each $s=0,1,2,\dots$, one introduces the quantities
$T_{n,s}(x):=\sum_{k=0}^{n}(k-nx)^{s}p_{n,k}(x),$
referred to as the algebraic moments of order $s$. In particular, we have $T_{n,0}(x)=1$ and $T_{n,1}(x)=0$, while higher-order moments can be expressed in terms of powers of $n$ and of the quantity X:=x(1x)X:=x(1-x).
Theorem 2.1 ([17]).
For a fixed $s=0,1,\dots$, $T_{n,s}(x)$ is a polynomial of degree $\leq s$ in $x$, and of degree $[s/2]$ in $n$. Moreover, if X:=x(1x),X:=x(1-x), $x\in[0,1]$, then
$\displaystyle T_{n,2s}(x)$ $=\sum_{j=1}^{s}a_{j,s}(X)n^{j}X^{j},$
$\displaystyle T_{n,2s+1}(x)$ $=(1-2x)\sum_{j=1}^{s}b_{j,s}(X)n^{j}X^{j},$
where $a_{j,s}$ and $b_{j,s}$ are polynomials of degree $\leq s-j,$ with coefficients independent of $n$.
From this theorem, it is possible to deduce several useful estimates. As a first example, for each $s=0,1,\dots$, there is a constant $A_{s}$ such that
$\displaystyle 0\leq T_{n,2s}(x)\leq A_{s}n^{s},$
$x\in[0,1]$. As a consequence, for each $\delta>0$ and $s=1,2,\dots$, there exists a constant $C=C(\delta,s)$ for which
$\displaystyle\sum_{|k/n-x|>\delta}p_{n,k}(x)\leq Cn^{-s},\hskip 8.53581ptn=1,2,\dots.$ (2)
For further details on these results and their applications, see Chapter 10 of [17].
@@ 3 Bernstein-type logarithmic operators
$\ln_{\mu}(x):=\ln(1+\mu+x),$ (3)
with $\mu>0$ and $x\in[0,1]$. Now, for $f\in C([0,1])$ and $n\in\mathbb{N}$ we define the new operator as
@@
$\displaystyle a_{n}(x):=\frac{\ln\left(1+\mu+\frac{x}{n}\right)-\ln\left(1+\mu\right)}{\ln\left(1+\mu+\frac{1}{n}\right)-\ln\left(1+\mu\right)},$ (5)
with $\mu>0$ and $x\in[0,1]$.
The function $a_{n}(x)$, that can be rewritten as
$\displaystyle a_{n}(x)=\frac{\ln\left(1+\frac{x}{n(1+\mu)}\right)}{\ln\left(1+\frac{1}{n(1+\mu)}\right)},$
is such that $a_{n}(0)=0$ and $a_{n}(1)=1$. Moreover, it is increasing and concave on $[0,1]$, and $a_{n}(x)\geq x$ for every $x\in[0,1]$ (see Figure 1).
Figure 1: Graphs of the functions $a_{n}(x)$ for $n=2,4,6$ compared with the identity function $e_{1}(x)=x$. The functions $a_{n}(x)$ approximate $e_{1}$ from above.
@@
Let $a_{n}(x)$ be defined as in (5). Then, the sequence $(a_{n}(x))_{n}$ converges uniformly to $e_{1}(x)\equiv x$ on $[0,1]$.
@@
There holds
$\displaystyle a_{n}^{\prime}(x)=\frac{1}{\ln\left(1+\frac{1}{n(1+\mu)}\right)}\frac{1}{1+\frac{x}{n(1+\mu)}}\frac{1}{n(1+\mu)},$
and so
$\displaystyle(a_{n}(x)-x)^{\prime}=\frac{1}{\ln\left(1+\frac{1}{n(1+\mu)}\right)}\frac{1}{1+\frac{x}{n(1+\mu)}}\frac{1}{n(1+\mu)}-1.$
Setting this expression equal to zero, we obtain
$\displaystyle 1+\frac{x}{n(1+\mu)}=\frac{1}{\ln\left(1+\frac{1}{n(1+\mu)}\right)}\frac{1}{n(1+\mu)},$
that is satisfied for
@@
$=$ $\frac{1-n(1+\mu)\ln{\left(1+\frac{1}{n(1+\mu)}\right)}}{\ln\left(1+\frac{1}{n(1+\mu)}\right)}.$
Since $(a_{n}(x)-x)^{\prime}>0$ if and only if $x<\bar{x}_{n}$, at the point $\bar{x}_{n}$ it is attained the maximum value of $(a_{n}(x)-x)$ in $[0,1]$, namely,
@@
$=\frac{\ln\left(1+\frac{1-n(1+\mu)\ln{\left(1+\frac{1}{n(1+\mu)}\right)}}{n(1+\mu)\ln\left(1+\frac{1}{n(1+\mu)}\right)}\right)}{\ln\left(1+\frac{1}{n(1+\mu)}\right)}-\frac{1-n(1+\mu)\ln{\left(1+\frac{1}{n(1+\mu)}\right)}}{\ln\left(1+\frac{1}{n(1+\mu)}\right)}$
@@
so that $\varepsilon_{n}\to 0^{+}$ as $n\to+\infty$, the limit $\lim_{n\to+\infty}[a_{n}(\bar{x}_{n})-\bar{x}_{n}]$ can be rewritten as
@@
To study the asymptotic behaviour, as $\varepsilon_{n}\to 0^{+}$, we apply the Taylor expansion of the logarithmic function and the geometric series expansion, obtaining
$\displaystyle\frac{\varepsilon_{n}}{\ln(1+\varepsilon_{n})}=\frac{1}{1-\frac{\varepsilon_{n}}{2}+o(\varepsilon_{n})}=1+\frac{\varepsilon_{n}}{2}+o(\varepsilon_{n}).$
Therefore,
@@
and so,
$\ln\left(\frac{\varepsilon_{n}}{\ln(1+\varepsilon_{n})}\right)-1=-1+\frac{\varepsilon_{n}}{2}-\frac{\varepsilon_{n}^{2}}{8}+o(\varepsilon_{n}^{2}).$
Moreover,
$\frac{1}{\ln(1+\varepsilon_{n})}=\frac{1}{\varepsilon_{n}-\frac{\varepsilon_{n}^{2}}{2}+o(\varepsilon_{n}^{2})},$
and so
$\displaystyle F(\varepsilon_{n})$ $=\frac{-1+\frac{\varepsilon_{n}}{2}-\frac{\varepsilon_{n}^{2}}{8}+o(\varepsilon_{n}^{2})}{\varepsilon_{n}-\frac{\varepsilon_{n}^{2}}{2}+o(\varepsilon_{n}^{2})}+\frac{1}{\varepsilon_{n}}$
$=\frac{o(\varepsilon_{n}^{2})}{\varepsilon_{n}^{2}-\frac{\varepsilon_{n}^{3}}{2}+o(\varepsilon_{n}^{3})}.$
Hence,
$\lim_{n\to+\infty}F(\varepsilon_{n})=0,$
and so the sequence $(a_{n}(x))_{n}$ converges uniformly to $e_{1}(x)\equiv x$ in $[0,1]$.
We observe that $(\mathscr{L}_{n})_{n}$ are positive linear operators that reproduce the logarithmic function $\ln_{\mu}(x)$, defined in (3). Indeed,
$\displaystyle\mathscr{L}_{n}(\ln_{\mu},x)$ $=\ln(1+\mu+x)\sum_{k=0}^{n}\ln\left(1+\mu+\frac{k}{n}\right)\frac{1}{\ln\left(1+\mu+\frac{k}{n}\right)}\hskip 2.84526ptp_{n,k}(a_{n}(x))$
$=\ln(1+\mu+x)\sum_{k=0}^{n}p_{n,k}(a_{n}(x))=\ln(1+\mu+x)=:\ln_{\mu}(x),$ (7)
since $\sum_{k=0}^{n}p_{n,k}(a_{n}(x))=1$, for every $x\in[0,1].$
The operators $(\mathscr{L}_{n})_{n}$ are connected to the classical Bernstein polynomials. In particular, if we denote
$f_{\mu}(x):=\frac{f(x)}{\ln(1+\mu+x)},\quad x\in[0,1],$ (8)
we find that
$\displaystyle\mathscr{L}_{n}f(x)$ $=\sum_{k=0}^{n}f\left(\frac{k}{n}\right)\frac{\ln(1+\mu+x)}{\ln\left(1+\mu+\frac{k}{n}\right)}\hskip 2.84526ptp_{n,k}(a_{n}(x))$
$=\ln(1+\mu+x)\sum_{k=0}^{n}f\left(\frac{k}{n}\right)\frac{1}{\ln_{\mu}(k/n)}p_{n,k}(a_{n}(x))$
$=\ln(1+\mu+x)\sum_{k=0}^{n}\frac{f}{\ln_{\mu}}\left(\frac{k}{n}\right)p_{n,k}(a_{n}(x))$
$=\ln(1+\mu+x)B_{n}\left(\frac{f}{\ln_{\mu}},a_{n}(x)\right)$
@@
where $B_{n}(f,x):=\sum_{k=0}^{n}f\left(\frac{k}{n}\right)p_{n,k}(x)$ is the classical Bernstein operator.
@@ 4 Pointwise and uniform convergence
In this section, we first provide a constructive proof of the pointwise and uniform convergence of the operators $(\mathscr{L}_{n})_{n}$.
@@
$\displaystyle I_{1}$ $=\left|\sum_{k=0}^{n}f\left(\frac{k}{n}\right)\frac{\ln_{\mu}(x)}{\ln_{\mu}(k/n)}\,p_{n,k}(a_{n}(x))-f(x)\sum_{k=0}^{n}\frac{\ln_{\mu}(x)}{\ln_{\mu}(k/n)}\,p_{n,k}(a_{n}(x))\right|$
$\leq\sum_{k=0}^{n}\Big|f\left(\frac{k}{n}\right)-f(x)\Big|\frac{\ln_{\mu}(x)}{\ln_{\mu}(k/n)}\,p_{n,k}(a_{n}(x))$
$\leq\Big\{\displaystyle\sum_{\left|x-\frac{k}{n}\right|\leq\delta}+\displaystyle\sum_{\left|x-\frac{k}{n}\right|>\delta}\Big\}\Big|f\left(\frac{k}{n}\right)-f(x)\Big|\frac{\ln_{\mu}(x)}{\ln_{\mu}(k/n)}\,p_{n,k}(a_{n}(x))$
@@
$=\|f\|_{\infty}\left|\sum_{k=0}^{n}\frac{\ln_{\mu}(x)}{\ln_{\mu}(k/n)}p_{n,k}(a_{n}(x))-1\right|$
$=\|f\|_{\infty}\left|\sum_{k=0}^{n}\left[\frac{\ln_{\mu}(x)}{\ln_{\mu}(k/n)}-\frac{\ln_{\mu}(k/n)}{\ln_{\mu}(k/n)}\right]p_{n,k}(a_{n}(x))\right|$
$\leq\|f\|_{\infty}\sum_{k=0}^{n}\frac{|\ln_{\mu}(x)-\ln_{\mu}(k/n)|}{\ln_{\mu}(k/n)}\,\,p_{n,k}(a_{n}(x))$
$\leq\|f\|_{\infty}\Big\{\sum_{\left|x-\frac{k}{n}\right|\leq\delta}+\sum_{\left|x-\frac{k}{n}\right|>\delta}\Big\}\frac{|\ln_{\mu}(x)-\ln_{\mu}(k/n)|}{\ln(1+\mu)}\,\,p_{n,k}(a_{n}(x))$
@@
when $\delta$ is the parameter for the uniform continuity of $\ln_{\mu}$ in $[0,1]$, correspondingly to $\varepsilon$. By similar reasonings as before, we have that
@@
We point out that the second part of the previous convergence result can be alternatively proved by means of the classical Korovkin theorem, which provides a sufficient condition to establish the uniform convergence of sequences of positive linear operators. Specifically, if $(T_{n})_{n\in\mathbb{N}}$ is a sequence of positive linear operators on $C([0,1])$ such that $T_{n}(e_{i})\to e_{i}$ uniformly on $[0,1]$ as $n\to+\infty$, with $e_{i}(t)=t^{i}$, $i=0,1,2$, then $T_{n}(f)\to f$ uniformly for every $f\in C([0,1])$.
This idea can be extended by introducing the concept of a Korovkin subset (see, e.g., [7, 6]), which allows for more general applications.
Definition 4.1.
Let $E$ and $F$ be Banach lattices, and let $T:E\to F$ be a positive linear operator. A subset $M\subset E$ is called a Korovkin subset for $T$ if, for every sequence $(T_{n})_{n\in\mathbb{N}}$ of positive linear operators from $E$ into $F$ satisfying:
  • (i)
    $\sup_{n\geq 1}\|T_{n}\|<+\infty$,
  • (ii)
    $\lim_{n\to\infty}T_{n}(g)=T(g)$ for all $g\in M$,
then
$\lim_{n\to\infty}T_{n}(f)=T(f)\quad\text{for all }f\in E.$
A useful test to verify whether a set is a Korovkin subset is the following.
Proposition 4.1 ([6]).
Let $M\subset C(X)$ and let $\mathcal{L}(M)$ be the linear span of $M$. Assume that for every pair $x\neq x_{0}$ in $X$, there exists $h\in\mathcal{L}(M)$, $h\geq 0$, such that $h(x)=0$ and $h(x_{0})>0$. Then $M$ is a Korovkin subset of $C(X)$.
For details and insights on Korovkin-type approximation theory, see, e.g. [15, 28, 7, 6] and the references therein.
Now, it is easy to see that certain logarithmic functions also form a Korovkin subset. In particular, we provide the following proposition.
@@
Given $\mu>0$ and $\lambda_{1},\lambda_{2}\in\mathbb{R}$, $0<\lambda_{1}<\lambda_{2},$ then $\{1,\ln_{\mu}^{\lambda_{1}},\ln_{\mu}^{\lambda_{2}}\}$ is a Korovkin subset of $C([0,1])$, where $\ln_{\mu}^{\lambda_{k}}(x):=\ln^{\lambda_{k}}(1+\mu+x)$ for every $x\in[0,1]$ and $k=1,2$.
@@
$\displaystyle h(x):=1+\frac{\lambda_{2}}{\lambda_{1}-\lambda_{2}}\frac{\ln^{\lambda_{1}}(1+\mu+x)}{\ln^{\lambda_{1}}(1+\mu+x_{0})}+\frac{\lambda_{1}}{\lambda_{2}-\lambda_{1}}\frac{\ln^{\lambda_{2}}(1+\mu+x)}{\ln^{\lambda_{2}}(1+\mu+x_{0})},$
@@
$\displaystyle h^{\prime}(x)$ $=\frac{\lambda_{1}\lambda_{2}}{\lambda_{1}-\lambda_{2}}\frac{1}{\ln^{\lambda_{1}}(1+\mu+x_{0})}\frac{\ln^{\lambda_{1}-1}(1+\mu+x)}{(1+\mu+x)}$
$\hskip 14.22636pt+\frac{\lambda_{1}\lambda_{2}}{\lambda_{2}-\lambda_{1}}\frac{1}{\ln^{\lambda_{2}}(1+\mu+x_{0})}\frac{\ln^{\lambda_{2}-1}(1+\mu+x)}{(1+\mu+x)}$
$\hskip-14.22636pt=\frac{\lambda_{1}\lambda_{2}}{\lambda_{2}-\lambda_{1}}\frac{\ln^{-1}(1+\mu+x)}{(1+\mu+x)}\frac{\ln^{\lambda_{1}}(1+\mu+x)}{\ln^{\lambda_{1}}(1+\mu+x_{0})}\left[-1+\frac{\ln^{\lambda_{2}-\lambda_{1}}(1+\mu+x)}{\ln^{\lambda_{2}-\lambda_{1}}(1+\mu+x_{0})}\right].$
@@
$\frac{\ln^{\lambda_{2}-\lambda_{1}}(1+\mu+x)}{\ln^{\lambda_{2}-\lambda_{1}}(1+\mu+x_{0})}>1,\quad\text{i.e.},\quad x>x_{0}.$
@@
Therefore, by Proposition 4.1, we conclude that $\{1,\ln_{\mu}^{\lambda_{1}},\ln_{\mu}^{\lambda_{2}}\}$ is a Korovkin subset of $C([0,1])$.
In the proof of Theorem 4.1, we saw that $(\mathscr{L}_{n}\ln^{2}_{\mu})_{n}$ converges uniformly to $\ln^{2}_{\mu}$ and, in a similar way it can be proved that $(\mathscr{L}_{n}1)_{n}$ converges uniformly to the constant function $1$ on $[0,1]$. Therefore, taking into account of (3) and the fact that, by Proposition 4.1, $\{1,\ln_{\mu},\ln^{2}_{\mu}\}$ is a Korovkin subset of $C([0,1])$, we can conclude that $(\mathscr{L}_{n}f)_{n}$ converges uniformly to $f$ on $[0,1]$, whenever $f\in C([0,1])$.

### v2

@@ Abstract
In this paper, we introduce a new class of positive linear operators that generalize the classical Bernstein operators. Specifically, we construct a sequence of operators that preserve the logarithmic function $\ln(1+\mu+x)$, with $\mu>0$ and $x\in[0,1]$. We prove pointwise and uniform convergence and we derive a quantitative estimate of the approximation error in terms of the modulus of continuity. We also obtain a Voronovskaja-type asymptotic formula, that is used to establish saturation results and inverse theorems. In particular, the saturation class of the considered approximation process is characterized by solving a second order differential equation. Shape-preserving properties, such as monotonicity, concavity and variation diminishing, are also investigated. Finally, a simple application to signal denoising is addressed.
@@ 1 Introduction
Over the years, many extensions of the Bernstein polynomials have been proposed in the literature, including, for example, the sampling-type operators ([14, 15]), the max-product operators ([16]), the neural network-type operators ([28]), the Szász-Mirakjan operators ([1, 2, 27, 10]), and many others ([26, 3, 34, 25, 24]).
@@
with $x\in[0,1]$. Such operators are naturally linked to the Bernstein polynomials $B_{n}$, since obviously
$\mathscr{G}_{n}(f,x)=\exp_{\mu}(x)\,B_{n}\left(\frac{f}{\exp_{\mu}},\tilde{a}_{n}(x)\right),$ (1)
where $\exp_{\mu}(x):=e^{\mu x}$. As a consequence, it can be shown that $\mathscr{G}_{n}$ preserves the exponential functions $\exp_{\mu}(x)$ and $\exp_{\mu}^{2}(x)$, i.e., for $x\in[0,1]$,
@@
Within this line of research, in the present paper we introduce a new class of positive linear operators, below denoted by $\mathscr{L}_{n}$, that preserve the logarithmic function $\ln_{\mu}(x)=\ln(1+\mu+x)$, with $\mu>0$ and $x\in[0,1]$. The idea is to replace the exponential weights in $\mathscr{G}_{n}$ with suitable logarithmic ones. Moreover, the functions $\tilde{a}_{n}$ are replaced by suitable concave approximants of the identity function involving $\ln_{\mu}(x)$. Differently from the exponential type operators, in the literature there is a lack of results concerning operators that preserve logarithmic functions. We mention that some results in this direction can be found in [33], but by means of integral operators of Mellin type, therefore in a completely different setting. Besides this, a further motivation for the introduction of the operators $\mathscr{L}_{n}$ is the fact that the properties of the logarithmic function could be useful to linearize nonlinear transformations of multiplicative type.
@@ 2 Bernstein-type logarithmic operators
$\ln_{\mu}(x):=\ln(1+\mu+x),\qquad x\in[0,1],$ (2)
where $\mu>0$ is a fixed parameter. Now, for a bounded function $f:[0,1]\rightarrow\mathbb{R}$ and $n\in\mathbb{N}$ we define the new operator as
@@
$\displaystyle a_{n}(x)\ :=\ \frac{\ln_{\mu}\left(\frac{x}{n}\right)-\ln\left(1+\mu\right)}{\ln_{\mu}\left(\frac{1}{n}\right)-\ln\left(1+\mu\right)}\ =\ \frac{\ln\left(1+\frac{x}{n(1+\mu)}\right)}{\ln\left(1+\frac{1}{n(1+\mu)}\right)},$ (4)
with $x\in[0,1]$. The function $a_{n}(x)$ is such that $a_{n}(0)=0$ and $a_{n}(1)=1$, is increasing and concave on $[0,1]$, and $a_{n}(x)\geq x$ for every $x\in[0,1]$.
@@
Let $a_{n}(x)$ be defined as in (4). Then, the sequence $(a_{n}(x))_{n}$ converges uniformly to the identity function $x$ on $[0,1]$.
@@
Since
$\displaystyle(a_{n}(x)-x)^{\prime}=\frac{1}{\ln\left(1+\frac{1}{n(1+\mu)}\right)}\frac{1}{1+\frac{x}{n(1+\mu)}}\frac{1}{n(1+\mu)}-1=0$
for
@@
$=$ $\frac{1-n(1+\mu)\ln{\left(1+\frac{1}{n(1+\mu)}\right)}}{\ln\left(1+\frac{1}{n(1+\mu)}\right)},$
and $(a_{n}(x)-x)^{\prime}>0$ if and only if $x<\bar{x}_{n}$, at the point $\bar{x}_{n}$ it is attained the maximum value of $(a_{n}(x)-x)$ in $[0,1]$, namely,
@@
there holds
@@
By the Taylor expansion of the logarithmic function and the geometric series,
@@
therefore
$\displaystyle F(\varepsilon_{n})\,$ $=\,\frac{-1+\frac{\varepsilon_{n}}{2}-\frac{\varepsilon_{n}^{2}}{8}+o(\varepsilon_{n}^{2})}{\varepsilon_{n}-\frac{\varepsilon_{n}^{2}}{2}+o(\varepsilon_{n}^{2})}+\frac{1}{\varepsilon_{n}}\,=\,\frac{o(\varepsilon_{n}^{2})}{\varepsilon_{n}^{2}-\frac{\varepsilon_{n}^{3}}{2}+o(\varepsilon_{n}^{3})}\to 0,\quad n\to+\infty,$
and so the sequence $(a_{n}(x))_{n}$ converges uniformly to $x$ in $[0,1]$.
We observe that $\mathscr{L}_{n}$ are positive linear operators that preserve the logarithmic function $\ln_{\mu}(x)$, defined in (2). Indeed,
$\displaystyle\mathscr{L}_{n}(\ln_{\mu},x)$ $=\ln_{\mu}(x)\sum_{k=0}^{n}\ln_{\mu}\left(\frac{k}{n}\right)\frac{1}{\ln_{\mu}\left(\frac{k}{n}\right)}\hskip 2.84526ptp_{n,k}(a_{n}(x))$
$=\ln_{\mu}(x)\sum_{k=0}^{n}p_{n,k}(a_{n}(x))=\ln_{\mu}(x),$ (6)
since $\sum_{k=0}^{n}p_{n,k}(a_{n}(x))=1$, for every $x\in[0,1].$ Nevertheless we point out that, differently from the exponential operators $\mathscr{G}_{n}$, $\mathscr{L}_{n}$ does not preserve the function $\ln_{\mu}^{2}(x)$.
The sequence $(\mathscr{L}_{n})_{n}$ is connected to the classical Bernstein polynomials. In particular, if we denote
$f_{\mu}(x):=\frac{f(x)}{\ln_{\mu}(x)},\qquad x\in[0,1],$ (7)
there holds
$\displaystyle\mathscr{L}_{n}f(x)$ $=\ln_{\mu}(x)\sum_{k=0}^{n}f\left(\frac{k}{n}\right)\frac{1}{\ln_{\mu}\left(\frac{k}{n}\right)}p_{n,k}(a_{n}(x))$
@@
which is the $\ln_{\mu}$-analogue of (1).
@@ 3 Pointwise and uniform convergence
In this section, we first provide a constructive proof of the pointwise and uniform convergence of the operators $\mathscr{L}_{n}$.
@@
$\displaystyle I_{1}$ $=\left|\sum_{k=0}^{n}f\left(\frac{k}{n}\right)\frac{\ln_{\mu}(x)}{\ln_{\mu}\left(\frac{k}{n}\right)}\,p_{n,k}(a_{n}(x))-f(x)\sum_{k=0}^{n}\frac{\ln_{\mu}(x)}{\ln_{\mu}\left(\frac{k}{n}\right)}\,p_{n,k}(a_{n}(x))\right|$
$\leq\sum_{k=0}^{n}\Big|f\left(\frac{k}{n}\right)-f(x)\Big|\frac{\ln_{\mu}(x)}{\ln_{\mu}\left(\frac{k}{n}\right)}\,p_{n,k}(a_{n}(x))$
$\leq\Big\{\displaystyle\sum_{\left|x-\frac{k}{n}\right|\leq\delta}+\displaystyle\sum_{\left|x-\frac{k}{n}\right|>\delta}\Big\}\Big|f\left(\frac{k}{n}\right)-f(x)\Big|\frac{\ln_{\mu}(x)}{\ln_{\mu}\left(\frac{k}{n}\right)}\,p_{n,k}(a_{n}(x))$
@@
$=\|f\|_{\infty}\left|\sum_{k=0}^{n}\frac{\ln_{\mu}(x)}{\ln_{\mu}\left(\frac{k}{n}\right)}p_{n,k}(a_{n}(x))-1\right|$
$=\|f\|_{\infty}\left|\sum_{k=0}^{n}\left[\frac{\ln_{\mu}(x)}{\ln_{\mu}\left(\frac{k}{n}\right)}-\frac{\ln_{\mu}\left(\frac{k}{n}\right)}{\ln_{\mu}\left(\frac{k}{n}\right)}\right]p_{n,k}(a_{n}(x))\right|$
$\leq\|f\|_{\infty}\sum_{k=0}^{n}\frac{|\ln_{\mu}(x)-\ln_{\mu}\left(\frac{k}{n}\right)|}{\ln_{\mu}\left(\frac{k}{n}\right)}\,\,p_{n,k}(a_{n}(x))$
$\leq\|f\|_{\infty}\Big\{\sum_{\left|x-\frac{k}{n}\right|\leq\delta}+\sum_{\left|x-\frac{k}{n}\right|>\delta}\Big\}\frac{|\ln_{\mu}(x)-\ln_{\mu}\left(\frac{k}{n}\right)|}{\ln(1+\mu)}\,\,p_{n,k}(a_{n}(x))$
@@
when $\delta$ is the parameter for the uniform continuity of $\ln_{\mu}(x)$ in $[0,1]$, correspondingly to $\varepsilon$. By similar reasonings as before, we have that
@@
We point out that the second part of the previous convergence result can be alternatively proved by means of the classical Korovkin theorem, which provides a sufficient condition to establish the uniform convergence of sequences of positive linear operators. Specifically, if $(T_{n})_{n\in\mathbb{N}}$ is a sequence of positive linear operators on $C([0,1])$ such that $T_{n}(e_{i})\to e_{i}$ uniformly on $[0,1]$ as $n\to+\infty$, with $e_{i}(x):=x^{i}$, $i=0,1,2$, then $T_{n}(f)\to f$ uniformly for every $f\in C([0,1])$.
This idea can be extended by recalling the well-known concept of Korovkin subset ([6, 5]). For details and insights on Korovkin-type approximation theory, see, e.g. [17, 31, 6, 5] and the references therein. We can prove the following.
@@
Given $\lambda_{1},\lambda_{2}\in\mathbb{R}$, $0<\lambda_{1}<\lambda_{2},$ then $\{1,\ln_{\mu}^{\lambda_{1}},\ln_{\mu}^{\lambda_{2}}\}$ is a Korovkin subset of $C([0,1])$.
@@
$\displaystyle h(x):=1+\frac{\lambda_{2}}{\lambda_{1}-\lambda_{2}}\frac{\ln_{\mu}^{\lambda_{1}}(x)}{\ln_{\mu}^{\lambda_{1}}(x_{0})}+\frac{\lambda_{1}}{\lambda_{2}-\lambda_{1}}\frac{\ln_{\mu}^{\lambda_{2}}(x)}{\ln_{\mu}^{\lambda_{2}}(x_{0})},$
@@
$\displaystyle h^{\prime}(x)$ $=\frac{\lambda_{1}\lambda_{2}}{\lambda_{1}-\lambda_{2}}\frac{1}{\ln_{\mu}^{\lambda_{1}}(x_{0})}\frac{\ln_{\mu}^{\lambda_{1}-1}(x)}{(1+\mu+x)}+\frac{\lambda_{1}\lambda_{2}}{\lambda_{2}-\lambda_{1}}\frac{1}{\ln_{\mu}^{\lambda_{2}}(x_{0})}\frac{\ln_{\mu}^{\lambda_{2}-1}(x)}{(1+\mu+x)}$
$\hskip-14.22636pt=\frac{\lambda_{1}\lambda_{2}}{\lambda_{2}-\lambda_{1}}\frac{\ln_{\mu}^{-1}(x)}{(1+\mu+x)}\frac{\ln_{\mu}^{\lambda_{1}}(x)}{\ln_{\mu}^{\lambda_{1}}(x_{0})}\left[-1+\frac{\ln_{\mu}^{\lambda_{2}-\lambda_{1}}(x)}{\ln_{\mu}^{\lambda_{2}-\lambda_{1}}(x_{0})}\right].$
@@
$\frac{\ln_{\mu}^{\lambda_{2}-\lambda_{1}}(x)}{\ln_{\mu}^{\lambda_{2}-\lambda_{1}}(x_{0})}>1,\quad\text{i.e.},\quad x>x_{0}.$
@@
Therefore, by Proposition 6.4 of [5], we conclude that $\{1,\ln_{\mu}^{\lambda_{1}},\ln_{\mu}^{\lambda_{2}}\}$ is a Korovkin subset of $C([0,1])$.
In the proof of Theorem 3.1, we saw that $(\mathscr{L}_{n}\ln^{2}_{\mu})_{n}$ converges uniformly to $\ln^{2}_{\mu}$ and, in a similar way it can be proved that $(\mathscr{L}_{n}1)_{n}$ converges uniformly to the constant function $1$ on $[0,1]$. Therefore, taking into account of (2) and the fact that, by Proposition 6.4 of [5], $\{1,\ln_{\mu},\ln^{2}_{\mu}\}$ is a Korovkin subset of $C([0,1])$, we can conclude that $(\mathscr{L}_{n}f)_{n}$ converges uniformly to $f$ on $[0,1]$, whenever $f\in C([0,1])$.